OpenStudy (anonymous):

Going against the current, a boat takes 6 hours to make a 120-mile trip. When the boat travels with the current on the return trip, it takes 5 hours. What is the rate of the boat in still water? 2 mph 20 mph 22 mph B??

OpenStudy (jack1):

this is the same q as before dude... just workin out a different variable

OpenStudy (anonymous):

im seeing if im correct lol DUDE

OpenStudy (jack1):

no, not B, what equation are u using to work it out? show ur working dude and we can help betta...?

OpenStudy (anonymous):

how is it not 20mph for 6 hours = 120 miles

OpenStudy (jack1):

What is the rate of the boat in STILL water?

OpenStudy (anonymous):

and c if it in still water correct

OpenStudy (jack1):

yep, c

OpenStudy (anonymous):

its 22 man...if you put x as the rate of the boat in still water and y as the rate of the current, then the combined rate with the current is x+y and against the current it is x−y you know the rate with the current is 120 5 =24 so x+y=24 and you know the rate against the current is 120 6 =20 so x−y=20 x+y=24 x−y=20 almost in your head but if you cannot come up with two numbers that add to 24 and whose difference is 20 then add up the two equations and get 2x=44 and so x=22